OpenStudy (anonymous):

Please please help D: I really have no idea how to put Part A in words. I know that the two line intersect at (1,4) but how do I answer A?? Here's the question: Explain why the x-coordinates of the points where the graphs of the equations y = 4^x and y = 2^x+2 intersect are the solutions of the equation 4^x=2^x+2.

OpenStudy (jtvatsim):

Man, I really hate these kind of questions. The assignments are totally missing the point... I'll think about what I'd say and let you know.

OpenStudy (anonymous):

Like I know if I substitute 1 for x and 4 for y then the equation will make sense, I just don't know how to answer the question

OpenStudy (jtvatsim):

"The x-coordinates of the points where the graphs of the equations intersect are the solutions because they satisfy the equation." Basically, maybe something extra like "A graph represents an equation, and we know that intersections mean the two equations give the same value at the x-coordinate, therefore they will make the equations equal and will be the solutions." I don't know... that's my best shot.

OpenStudy (nincompoop):

how do you "know" that those are the intersections?

OpenStudy (anonymous):

Thank you @jtvatsim and @nincompoop I used Desmos to graph both of the lines and they intersected at (1,4) :)

OpenStudy (nincompoop):

oh wow, so you didn't actually solve it no wonder you can't put it into words

OpenStudy (anonymous):

What is there to solve, all I needed was the point of intersection to get the answer to the problem. And just for your knowledge, I answered the question a few seconds before @jtvatsim replied. Here is what I answered if you want to know so badly -__-: They are the solutions to the equation 4^x=2^x+2 because (1,4) is the point of intersection and when we plug 1 for 'x' and 4 for 'y' we do 4=4^1 we get 4=4. And when we do 4=2^1+2 we get 4=4 Yes. So, 4^x is the same as 2^x+2 because 'x' equals 1 and 'y' equals 4. Bye